OpenStudy (anonymous):

For the simple harmonic motion equation, d = 5sin(2 pi t), what is the period? If necessary, use the slash (/) to represent a fraction. ***My Answer:1*** Is this right?

OpenStudy (anonymous):

\[period =\frac{ 2\pi }{ b}\] What would the 'b' Value be?

OpenStudy (anonymous):

You're right, the answer is 1. \(\large \frac{2\pi}{2\pi}=1\)

OpenStudy (anonymous):

Oh ok! So the b value is 2pi! Thank you!